OpenStudy (anonymous):

how do I simplify the radical expression root 12 over root 3

OpenStudy (anonymous):

you can combine the radicals into one:\[\sqrt{\frac{ 12 }{ 3 }}\]

OpenStudy (anonymous):

and I think you know what 12 divided by 3 is!

OpenStudy (tkhunny):

Find the perfect squares and pull them out. \(\sqrt{12} = \sqrt{4\cdot 3} = \sqrt{4}\cdot \sqrt{3} = 2\cdot\sqrt{3}\)

OpenStudy (anonymous):

ok, the answer that I got is 2.

OpenStudy (anonymous):

Am I correct?

OpenStudy (anonymous):

bingo!

OpenStudy (tkhunny):

Why do you doubt?

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

i just wanted to make sure.

OpenStudy (tkhunny):

Well, stop it. :-) Be sure in your mind. Do 10 more such problems and earn some confidence. Good work.

OpenStudy (anonymous):

okay, thank you